OpenStudy (dmcdowell):

please help im i gotta get this question asap A skyscraper is 396 meters tall. At a certain time of day, it casts a shadow that is 332 meters long. At what angle is the sun above the horizon at that time? 57° 50° 40° 33°

OpenStudy (anonymous):

\[\tan(\theta)=\frac{369}{332}\]solve for \(\theta\)

OpenStudy (dmcdowell):

ok one sec @satellite73

OpenStudy (dmcdowell):

i got it thanks